Project

General

Profile

NDOS Commissioning Meeting 2011/03/24

Attendance: Jon, Susan, Enrique, Gavin, Carl B., Minerba, Mat, Mark Mathis, Chad, Denis, Gary, Athans, Peter, Ryan, Leon, Jarek,

Agenda:

  1. DCS Status & Plans
  2. Discussion of new logbook (ECL)
  3. Installation update
  4. DSO Studies
  5. DAQ/Electronics Update & Plans
  6. AOB

DCS:

  • Chad: took some effort to find Pb-free solder (got it from Granger), will begin wiring up RTDs early next week. RTDs are very small, so they'll be covered up with colored tape. Most RTD readouts will be installed in rack 2. May need some help next week. Also working on making monitoring program
  • AC in NDOS building was broken. Bill Moorehouse is the building manager.
  • Athans tried to read the temperature register of some FEBs on the teststand and got only 0s. He asked who he should contact; Peter suggested Qiming, but in general questions like that should simply be sent to the DAQ email list. Athans will follow up with an email after the meeting.

ECL Discussion

  • No showstoppers, but it is greatly desired to have a search facility which understands "wildcards"
  • Jon will inform the dB Apps group about the decision to move forward with ECL. The plan is as follows:
    1. On Sunday, we will delete ECL entries made on 3/21/2011 or after and then re-import all CRL entries from 3/21/2011 on.
    2. Monday by 8:00: on dbweb1 and dbweb2 we will replace .../index.jsp with HTML notifying the user about new URL: http://dbweb0.fnal.gov/ECL/nova this will effective close NOvA CRL and redirect users to ECL.
    3. Monday by 9:00: copy all entries created since Sunday night, keeping their entry ids.

Installation/Filling

  • Minerba/Mat: filling of verticals is done, filling shifts will resume tomorrow afternoon. Looking for volunteers to continue filling on Saturday and early next week.
  • Denis: yesterday, 9 cap. boards were installed, one was unsuccessful (wire was cut by accident during installation). DSO scans show an improvement; most channels show no noticeable noise levels; one showed a lot of noise. Tightening grounding wires seemed to fix the worst offender.
  • Schedule for remaining cap board (Leon): Stuart says that boards are on the way and parts are in hand. Need to ship these off to the assembler (late next week). So boards will be available to install in ~2 weeks.

DAQ Software:

  • Peter:
    • moved forward again on new linux dist. on DCMs. Still seems ok, but still some issues with basic processes crashing (eg, top, ps)
    • DDS message buffer was increased and greatly reduced the number of lost trigger messages. Today lost trigger messages were observed around the time a new subrun message is sent out from RC.

AOB

  • Andrew: USB port for DCMs at NDOS just came in. Requested time on detector to install. The plan is to run cables tomorrow morning for a couple of DCMs and hook a few DCMs up. If all goes well, work will be completed next week.
  • Susan:
    • Noted that the DAQ is much more stable than the last time she was on shift (just a few weeks ago)
    • Asked if we'd like results from DataCheck shown weekly at this meeting. An emphatic YES! Plots will be shown next Thursday.
    • Noted that the OnMon is janky Mark Messier is on shift next week and will hopefully have time to look into it.